# Clustering by Denoising: Latent plug-and-play diffusion for single-cell data

### v2

@@ 1 Introduction
We frame single-cell denoising as an inverse problem: recovering clean gene expression from noisy measurements without imposing strong generative assumptions. Our method leverages the high-quality signal from reference datasets (e.g., SMART-seq2, Picelli et al. 2013; 2014) to learn a powerful denoising prior, which is then applied to enhance noisier target data (e.g., droplet-based scRNA-seq, Klein et al. 2015; Macosko et al. 2015) for improved clustering. We operationalize this under the plug-and-play (PnP) paradigm (Venkatakrishnan et al., 2013a; Zhang et al., 2021; Chan et al., 2016; Ryu et al., 2019), which integrates powerful denoising priors with measurement models. Mainstream PnP diffusion frameworks (Zhu et al., 2023; Go et al., 2023; Wu et al., 2024; Coeurdoux et al., 2024; Xu & Chi, 2024) enables this through combining likelihoods via iterative refinement (e.g., Gibbs sampling), where each denoising step is followed by controlled noise reintroduction to enforce data consistency. This principled approach enables us to denoise beyond test data quality by transferring patterns from high-signal reference data to noisier technologies, even under mild distribution shifts between the reference and target distribution. However, directly applying image-based PnP frameworks to single-cell data is challenging. Unlike images where pixel noise is largely independent, gene expression data exhibits intrinsic low-rank structure and complex correlations. Moreover, denoising must preserve relational structure between cells for accurate clustering and annotation. Standard dimensionality reduction (e.g., PCA) can collapse distinct cell types, making it impossible to guide denoising accurately.
@@
Our framework diverges from existing Bayesian approaches for single-cell analysis by removing their need for restrictive generative modeling. While variational autoencoders (VAEs) (Lopez et al., 2018a; Gayoso et al., 2019; Grønbech et al., 2020) are difficult to train and rely on strong likelihood assumptions, more recent approximate message passing methods with empirical Bayes denoisers (Zhong et al., 2022; Nandy & Ma, 2024) still require parametric noise modeling, operate purely in the latent space, and scale poorly to high-dimensional latent spaces. In contrast, our approach requires no explicit generative model or pre-processing for noise structure, instead learning it directly from data.
@@
Our experimental results demonstrate consistent performance under various mis-specifications of the data-generating process in synthetic settings (Section 4). In real-world single-cell experiments, our method shows strong potential for leveraging high-signal training data to improve denoising in low-signal datasets and to denoise beyond the training distribution by averaging (Section 5).
@@ 2 Single-Cell Atlas Construction via Posterior Sampling
    Unsupervised Clustering with Post-Hoc Evaluation: Denoised embeddings $\{{\widehat{U}_{i}}\}_{i=1}^{n}$ are clustered without using label information to obtain $K$ clusters. The quality of these clusters is evaluated by comparing them to the known cell type labels $L_{i}\in{1,\ldots,C}$ using the metrics adjusted rand index (ARI) (Hubert & Arabie, 1985), average silhouette score (Rousseeuw, 1987), normalized mutual information (NMI) (Strehl & Ghosh, 2002), and cell type locally invariant Simpson’s index (cLISI) Korsunsky et al. (2019) [^1] For detailed description of these metrics, see Section D., which quantify the agreement between the data-driven clusters and biological annotations.
Our framework is fully unsupervised, relying only on expression data. [^2] While labels are not used in denoising, we explore label-augmented diffusion training in Section 5. We benchmark performance on synthetic data with known ground truth and on real datasets, noting that real-world labels may be imperfect (and thus annotation accuracy thus provides an indirect measure of atlas quality).
@@ 3 Method
An illustration of the denoising can be found in Figure A.1. This approach preserves the flexibility of diffusion priors while maintaining tractable posterior inference, providing a scalable and uncertainty-aware framework for single-cell atlas construction.
@@
    Likelihood alignment (Line 4) is implemented using either a general proximal scheme (Xu & Chi, 2024) or a closed-form Gaussian update when $f$ is Gaussian (Proposition 3.1). Unlike the likelihood alignment step in Eq. (5), Line 4 operates in the original $d$-dimensional data space, reintroducing noise through the likelihood function $\log f(X_{q}-\widehat{V}Z_{q})$.
@@ 4 Evaluation on synthetic data
    Setup 1 (matched train–test distribution). We generate $\mathcal{D}^{{(t)}}$ from the same DGP as $\mathcal{D}^{{(r)}}$ to isolate posterior denoising effects. This evaluates whether DICE improves population separation in latent space versus PCA, mirroring standard atlas workflows where query cells are mapped via denoised embeddings (Nandy & Ma, 2024).
@@
Uncertainty in the embeddings   We visualize our confidence set construction for two fixed points: (i) the center of Cluster 2 and (ii) the midpoint between the two cluster centers. For each input, we run DICE 500 times and project the resulting embeddings onto the same UMAP as the training data. When the input lies at the center of a cluster (Figure 2, left), all embeddings map consistently to Cluster 2, indicating high confidence in the cluster assignment. In contrast, when the input lies between clusters (Figure 2, right), the embeddings are split across both clusters, reflecting uncertainty in the assignment. Such uncertainty could be advantageous for downstream tasks designed to incorporate soft labels. Additional plots in subsection E.1 show how the parameter $\rho$ directly controls the size of the confidence sets. Thus, $\rho$ must be tuned to balance coverage and performance.
@@
Figure 2: UMAP visualizations of 500 runs of DICE on the same input point in Setup 1 and $\rho=0.1$. Training data are shown in grey. Left: center of Cluster 2; Right: midpoint between Clusters 1 and 2.
@@ 5 Evaluation on single-cell data
We evaluate the effectiveness of DICE in denoising single-cell gene expression profiles using two publicly available single-cell RNA-seq datasets: the CITE-seq dataset from Hao et al. (2021) and the human fetal brain development datasets from Polioudakis et al. (2019) and Nowakowski et al. (2017). These datasets originate from distinct tissues and capture diverse cellular populations. They also differ in their relative signal strengths, allowing us to examine the ability of our method to handle both complex cell distributions and varying signal-to-noise regimes.
Diffusion Training   For each dataset, we first select the latent dimension $k$ using the elbow of the singular-value spectrum, and then project the training data onto this $k$-dimensional space via PCA. This yields the training embeddings $\{\widehat{U}_{i}\}_{i=1}^{m}$ and the loading matrix $\widehat{V}$ used by DICE. We train the same diffusion architecture across datasets using AdamW with a cosine-annealed learning-rate schedule. Additional details on the training pipeline, including precise model architecture, are provided in Section C of the appendix.
@@ 5.1 Analysis of the CITE-seq dataset
Dataset   CITE-seq dataset (Hao et al., 2021) consists of RNA profiles for 20,729 genes across 152,094 PBMCs with paired antibody-derived tags (ADT) measurements. We focus solely on the transcriptomic modality and uniformly subsample 10,000 cells for analysis. Ground-truth labels are provided, and we adopt the L2-level granularity, which distinguishes $\sim$30 immune subtypes.
Preprocessing and denoising   We applied standard QC and pre-processing to the raw UMI counts data (see details in Section H), and considered the top 3,000 highly variable genes (Seurat v3 criterion) to construct the data matrix $X\in\mathbb{R}^{10{,}000\times 3{,}000}$. We split into 9,000 training and 1,000 held-out test cells with preserved label proportions. Similar to Setup 1, we test DICE’s ability in denoising beyond the quality of the training dataset in this setting. The training set was used to fit the diffusion model, and the test set was reserved for evaluation.
@@
Evaluation   We compare UMAP (McInnes et al., 2020) visualizations of the 1,000 held-out test cells denoised by DICE against their PCA projections without denoising (Figure 3). Denoised embeddings show clearer segregation of immune subtypes, with marked improvement in separating sub lineages of CD4 and CD8 T cells. The DICE atlas also performs better in seggregating the MAIT cells from the other T cells. These populations are notoriously difficult to resolve from RNA alone (Hao et al., 2021), a challenge that originally motivated multimodal approaches combining RNA and protein. In the unimodal setting considered here, however, popular toolkits such as Seurat rely directly on PCA embeddings. Our results indicate that denoising PCA embeddings with priors learned from high-signal training data substantially improves cluster separation, supporting more reliable annotation of held-out cells. We additionally benchmarked the DICE embeddings against several widely used denoising procedures for RNA-seq analysis—MAGIC (van Dijk et al., 2018a), ALRA (Linderman et al., 2022a), $k$-nearest-neighbor smoothing (used in Seurat v3 (Stuart et al., 2019)), and non-negative matrix factorization (NMF)–based embeddings [^3] See Section I for details on the benchmarking procedures and their implementation., using the clustering metrics ARI, NMI, cLISI, and V-measure described in Section D. The results, summarized in Table 2, show that DICE consistently outperforms these popular denoising pipelines across most clustering metrics.
@@ 5.2 Analysis of Human Fetal Brain Development Datasets
Datasets   We evaluate the effectiveness of DICE in transferring cell type information learned from high-signal training data to related low-signal test data using scRNA-seq datasets from Nowakowski et al. (2017) and Polioudakis et al. (2019). Both datasets profile human fetal brain tissue during development. The Nowakowski et al. (2017) dataset includes cells from primary cortical and medial ganglionic eminence (MGE) samples across multiple stages spanning peak neurogenesis, whereas the Polioudakis et al. (2019) dataset focuses on cells from the neocortex during mid-gestation (15–21 post-conception weeks). While the cell types profiled in the two datasets are related, they are not identical due to differences in the sampled tissue. This contrasts with the CITE-seq dataset considered previously, where training and test cells originated from the same source. This experiment evaluates the robustness of DICE to realistic distributional changes that occur during cross-dataset cell-type label transfer.
@@
Evaluation   We compare UMAP visualizations for the embeddings constructed from the test data using DICE and the embeddings obtained via the PCA transform learned from the training data. Denoising yielded substantially improved biological interpretability, producing more coherent and homogeneous clusters of similar cell types. A notable improvement is found in the embeddings of the cells belonging to the canonical excitatory trajectory $\mathrm{RG}\!\to\!\mathrm{IPC}\!\to\!\mathrm{nEN}\!\to\!\mathrm{EN}.$ This trajectory is visually continuous and easy to follow in the DICE embedding, whereas it appears fragmented and noisy under PCA. Similar to the CITE-seq example, we benchmarked the DICE embeddings against MAGIC, ALRA, $k$-nearest smoothing, and NMF, using the same set of clustering metrics. The results summarized in 2, show that DICE consistently outperforms all competing denoising methods across all four metrics, validating our claim that diffusion-based denoising sharpens lineage relationships and resolves biologically relevant subpopulations better than existing techniques in the literature.
@@ Appendix A Diffusion Training
We implemented all experiments in Python. Neural network training was performed in PyTorch (Paszke et al., 2019) on GPU-accelerated hardware, using a mix of AWS g4dn.xlarge EC2 instances and a dedicated Ubuntu server with an NVIDIA RTX PRO 6000 GPU.
@@
Inference proceeds via Gibbs sampling, alternating between the prior and likelihood steps. The prior step, which involves sampling from the diffusion model, is the most computationally expensive. As $\rho$ determines the starting timestep of the diffusion model $t$, the duration of the prior sampling is dependent on $\rho$, larger values of $\rho$ result in longer runtime. The of total runtime for denoising the test set of the CITE-seq dataset with 100 Gibbs iterations and $\rho$ ranging from $5.0$ to $0.5$ took 12 minutes.
@@ Appendix D Discussion on the Clustering Metrics
Figure A.1: Visual sketch of the DICE denoising procedure. Training stage: (1) The factor loading matrix $\hat{V}$ is estimated on the reference dataset. (2) A diffusion model is trained to capture the underlying biological variation. Inference stage: (3) The initial latent representation $U^{(0)}$ is obtained using $\hat{V}$. Iteratively: (4) Noise is added to $U^{(s)}$ to produce $Z^{(s)}$. (5) $Z^{(s)}$ is denoised using the learned diffusion model, yielding $U^{(s+w)}$.
@@ H.2 Human Neo-Cortex Datasets
Because the two studies differ in experimental protocols and developmental stages, only 17,638 genes were shared across both datasets. From these, we selected the top 5,000 highly variable genes (HVGs) within each dataset and retained the intersection of these HVG sets, yielding 785 shared highly variable genes. This resulted in the matrices $X_{\mathrm{now}}\in\mathbb{R}^{3495\times 785}$ and $X_{\mathrm{pol}}\in\mathbb{R}^{15126\times 785}$.
@@ Appendix I Implementation Details for Benchmarking
  • Non-negative Matrix Factorization (NMF) is a commonly used alternative to PCA for denoising and latent factor extraction in scRNA-seq analysis (Kotliar et al., 2019; Carbonetto et al., 2021). We used the NMF implementation from scikit-learn (Pedregosa et al., 2011), selecting the latent dimension to match the number of PCs used for the DICE embeddings.

### v3

@@ 1 Introduction
We frame single-cell denoising as an inverse problem: recovering clean gene expression from noisy measurements without imposing strong generative assumptions. Our method leverages the high-quality signal from reference datasets (e.g., SMART-seq2, Picelli et al. 2013; 2014) to learn a powerful denoising prior, which is then applied to enhance noisier target data (e.g., droplet-based scRNA-seq, Klein et al. 2015; Macosko et al. 2015) for improved clustering. We operationalize this under the plug-and-play (PnP) paradigm (Venkatakrishnan et al., 2013a; Zhang et al., 2021; Chan et al., 2016; Ryu et al., 2019), which integrates powerful denoising priors with measurement models. Mainstream PnP diffusion frameworks (Zhu et al., 2023; Go et al., 2023; Wu et al., 2024; Coeurdoux et al., 2024; Xu & Chi, 2024) enable this through combining likelihoods via iterative refinement (e.g., Gibbs sampling), where each denoising step is followed by controlled noise reintroduction to enforce data consistency. This principled approach enables us to denoise beyond test data quality by transferring patterns from high-signal reference data to noisier technologies, even under mild distribution shifts between the reference and target distribution. However, directly applying image-based PnP frameworks to single-cell data is challenging. Unlike images where pixel noise is largely independent, gene expression data exhibits intrinsic low-rank structure and complex correlations. Moreover, denoising must preserve relational structure between cells for accurate clustering and annotation. Standard dimensionality reduction (e.g., PCA) can collapse distinct cell types, making it impossible to guide denoising accurately.
@@
Our framework diverges from existing Bayesian approaches for single-cell analysis by removing their need for restrictive generative modeling. While variational autoencoders (VAEs) (Lopez et al., 2018a; Gayoso et al., 2021; Grønbech et al., 2020) are difficult to train and rely on strong likelihood assumptions, more recent approximate message passing methods with empirical Bayes denoisers (Zhong et al., 2022; Nandy & Ma, 2024) still require parametric noise modeling, operate purely in the latent space, and scale poorly to high-dimensional latent spaces. In contrast, our approach requires no explicit generative model or pre-processing for noise structure, instead learning it directly from data.
@@
Our experimental results demonstrate consistent performance under various mis-specifications of the data-generating process in synthetic settings (Section 4). In real-world single-cell experiments, our method shows strong potential for leveraging high-signal training data to improve denoising in low-signal datasets and to denoise beyond the training distribution by averaging (Section 5). The source code of DICE is publicly available. [^1] https://github.com/dommeier/dice
@@ 2 Single-Cell Atlas Construction via Posterior Sampling
    Unsupervised Clustering with Post-Hoc Evaluation: Denoised embeddings $\{{\widehat{U}_{i}}\}_{i=1}^{n}$ are clustered without using label information to obtain $K$ clusters. The quality of these clusters is evaluated by comparing them to the known cell type labels $L_{i}\in{1,\ldots,C}$ using the metrics adjusted rand index (ARI) (Hubert & Arabie, 1985), average silhouette score (Rousseeuw, 1987), normalized mutual information (NMI) (Strehl & Ghosh, 2002), and cell type locally invariant Simpson’s index (cLISI) Korsunsky et al. (2019), [^2] For detailed description of these metrics, see Appendix D. which quantify the agreement between the data-driven clusters and biological annotations.
Our framework is fully unsupervised, relying only on expression data. [^3] While labels are not used in denoising, we explore label-augmented diffusion training in Section 5. We benchmark performance on synthetic data with known ground truth and on real datasets, noting that real-world labels may be imperfect (and thus annotation accuracy provides an indirect measure of atlas quality).
@@ 3 Method
An illustration of the method can be found in Figure A.1. This approach preserves the flexibility of diffusion priors while maintaining tractable posterior inference, providing a scalable and uncertainty-aware framework for single-cell atlas construction.
@@
    Likelihood alignment (Line 4) is implemented using either a general proximal scheme (Xu & Chi, 2024) or a closed-form Gaussian update when $f$ is Gaussian (Proposition 3.1). Unlike the prior alignment step in Eq. (5), Line 4 operates in the original $d$-dimensional data space, reintroducing noise through the likelihood function $\log f(X_{q}-\widehat{V}Z_{q})$.
@@ 4 Evaluation on synthetic data
    Setup 1 (matched train–test distribution). We generate $\mathcal{D}^{{(t)}}$ from the same generative distribution as $\mathcal{D}^{{(r)}}$ to isolate posterior denoising effects. This evaluates whether DICE improves population separation in latent space versus PCA, mirroring standard atlas workflows where query cells are mapped via denoised embeddings (Nandy & Ma, 2024).
@@
Uncertainty in the embeddings   We visualize our confidence set construction for two fixed points: (i) the center of Cluster 2 and (ii) the midpoint between the two cluster centers. For each input, we run DICE 500 times and project the resulting embeddings onto the same UMAP as the training data. When the input lies at the center of a cluster (Figure 2, left), all embeddings map consistently to Cluster 2, indicating high confidence in the cluster assignment. In contrast, when the input lies between clusters (Figure 2, right), the embeddings are split across both clusters, reflecting uncertainty in the assignment. Such uncertainty could be advantageous for downstream tasks designed to incorporate soft labels. Additional plots in Appendix E show how the parameter $\rho$ directly controls the size of the confidence sets. Thus, $\rho$ must be tuned to balance coverage and performance.
@@
Figure 2: UMAP visualizations of 500 runs of DICE on the same input point in Setup 1 and $\rho=0.1$. Training data are shown in grey. Left: center of Cluster 2; Right: midpoint between Clusters 1 & 2.
@@ 5 Evaluation on single-cell data
We evaluate the effectiveness of DICE in denoising gene expression profiles using two publicly available single-cell RNA-seq datasets: the CITE-seq dataset from Hao et al. (2021) and the human fetal brain development datasets from Polioudakis et al. (2019) and Nowakowski et al. (2017) (see Appendix G for dataset provenance). These datasets originate from distinct tissues and capture diverse cellular populations. They also differ in their relative signal strengths, allowing us to examine the ability of DICE to handle both complex cell distributions and varying signal-to-noise regimes.
Diffusion Training   For each dataset, we first select the latent dimension $k$ using the elbow of the singular-value spectrum, and then project the training data onto this $k$-dimensional space via PCA. This yields the training embeddings $\{\widehat{U}_{i}\}_{i=1}^{m}$ and the loading matrix $\widehat{V}$ used by DICE. We train the same diffusion architecture across datasets using AdamW with a cosine-annealed learning-rate schedule. Additional details on the training pipeline, including precise model architecture, are provided in Appendix C. We evaluate the influence of picking the PCA dimension ($k$) in Appendix F.
@@ 5.1 Analysis of the CITE-seq dataset
Dataset   The CITE-seq dataset (Hao et al., 2021) consists of RNA profiles for 20,729 genes across 152,094 PBMCs with paired antibody-derived tags (ADT) measurements. We focus solely on the transcriptomic modality and uniformly subsample 10,000 cells for analysis. Ground-truth labels are provided, and we adopt the L2-level granularity, which distinguishes $\sim$30 immune subtypes.
Preprocessing and denoising   We applied standard QC and pre-processing to the raw UMI counts data (see details in Appendix H), and considered the top 3,000 highly variable genes (Seurat v3 criterion) to construct the data matrix $X\in\mathbb{R}^{10{,}000\times 3{,}000}$. We split into 9,000 training and 1,000 held-out test cells with preserved label proportions. Similar to Setup 1, we test DICE’s ability in denoising beyond the quality of the training dataset in this setting. The training set was used to fit the diffusion model, and the test set was reserved for evaluation.
@@
Evaluation   We compare UMAP (McInnes et al., 2018) visualizations of the 1,000 held-out test cells denoised by DICE against their PCA projections without denoising (Figure 3). Denoised embeddings show clearer segregation of immune subtypes, with marked improvement in separating sublineages of CD4 and CD8 T cells. The DICE atlas also performs better in segregating the MAIT cells from the other T cells. These populations are notoriously difficult to resolve from RNA alone (Hao et al., 2021), a challenge that originally motivated multimodal approaches combining RNA and protein. In the unimodal setting considered here, however, popular toolkits such as Seurat rely directly on PCA embeddings. Our results indicate that denoising PCA embeddings with priors learned from high-signal training data substantially improves cluster separation, supporting more reliable annotation of held-out cells. We additionally benchmarked the DICE embeddings against several widely used denoising procedures for RNA-seq analysis—MAGIC (van Dijk et al., 2018a), ALRA (Linderman et al., 2022a), $k$-nearest-neighbor smoothing (used in Seurat v3 (Stuart et al., 2019)), and non-negative matrix factorization (NMF)–based embeddings, [^4] See Appendix I for details on the benchmarking procedures and their implementation. using the clustering metrics ARI, NMI, cLISI, and V-measure described in Section D. The results (Table 2) show that DICE consistently outperforms these popular denoising pipelines across most clustering metrics.
@@ 5.2 Analysis of Human Fetal Brain Development Datasets
Datasets   We evaluate the effectiveness of DICE in transferring cell type information learned from high-signal training data to related low-signal test data using scRNA-seq datasets from Nowakowski et al. (2017) and Polioudakis et al. (2019). Both datasets profile human fetal brain tissue during development. The Nowakowski et al. (2017) dataset includes cells from primary cortical and medial ganglionic eminence (MGE) samples across multiple development stages, whereas the Polioudakis et al. (2019) dataset focuses on cells from the neocortex during mid-gestation. While the cell types profiled in the two datasets are related, they are not identical due to differences in the sampled tissue. This contrasts with the CITE-seq dataset considered previously, where training and test cells originated from the same source. This experiment evaluates the robustness of DICE to realistic distributional changes that occur during cross-dataset cell-type label transfer.
@@
Evaluation   We compare UMAP visualizations for the embeddings constructed from the test data using DICE and the embeddings obtained via the PCA transform learned from the training data. Denoising yielded substantially improved biological interpretability, producing more coherent and homogeneous clusters of similar cell types. A notable improvement is found in the embeddings of the cells belonging to the canonical excitatory trajectory $\mathrm{RG}\!\to\!\mathrm{IPC}\!\to\!\mathrm{nEN}\!\to\!\mathrm{EN}.$ This trajectory is visually continuous and easy to follow in the DICE embedding, whereas it appears fragmented and noisy under PCA. Similar to the CITE-seq example, we benchmarked the DICE embeddings against MAGIC, ALRA, $k$-nearest smoothing, and NMF, using the same set of clustering metrics. The results summarized in 2, show that DICE consistently outperforms all competing denoising methods across all four metrics, validating our claim that diffusion-based denoising sharpens lineage relationships and resolves biologically relevant subpopulations better than existing techniques.
@@ Appendix A Diffusion Training
We implemented all experiments in Python. Neural network training was performed in PyTorch (Paszke et al., 2019) on GPU-accelerated hardware, using a mix of AWS g4dn.xlarge EC2 instances and a dedicated Ubuntu server with an NVIDIA RTX PRO 6000 GPU. Our whole code is available online at
@@
Inference proceeds via Gibbs sampling, alternating between the prior and likelihood steps. The prior step, which involves sampling from the diffusion model, is the most computationally expensive. As $\rho$ determines the starting timestep of the diffusion model $t$, the duration of the prior sampling is dependent on $\rho$, larger values of $\rho$ result in longer runtime. The total runtime for denoising the test set of the CITE-seq dataset with 100 Gibbs iterations and $\rho$ ranging from $5.0$ to $0.5$ took 12 minutes.
@@ Appendix D Discussion on the Clustering Metrics
Figure A.1: Visual sketch of the DICE denoising procedure. Training stage: (1) The factor loading matrix $\hat{V}$ is estimated on the reference dataset. (2) A diffusion model is trained to capture the underlying biological variation. Inference stage: (3) The initial latent representation $U^{(0)}$ is obtained using $\hat{V}$. Iteratively: (4) Noise is added to $U^{(s)}$ to produce $Z^{(s)}$. (5) $Z^{(s)}$ is denoised using the learned diffusion model, yielding $U^{(s+1)}$.
@@ H.2 Human Neo-Cortex Datasets
Because the two studies differed in experimental protocols and developmental stages, only 17,638 genes were shared across both datasets. From these, we selected the top 5,000 highly variable genes (HVGs) within each dataset and retained the intersection of these HVG sets, yielding 785 shared highly variable genes. This resulted in the matrices $X_{\mathrm{now}}\in\mathbb{R}^{3495\times 785}$ and $X_{\mathrm{pol}}\in\mathbb{R}^{15126\times 785}$.
@@ Appendix I Implementation Details for Benchmarking
  • Non-negative Matrix Factorization (NMF) is a commonly used alternative to PCA for denoising and latent factor extraction in scRNA-seq analysis (Kotliar et al., 2019; Carbonetto et al., 2026). We used the NMF implementation from scikit-learn (Pedregosa et al., 2011), selecting the latent dimension to match the number of PCs used for the DICE embeddings.